# Yet Another Representation of Binary Decision Trees: A Mathematical Demonstration

### v7

@@ 3.1.3 The Bitvector Matrix
The bitvector matrix $B$ in (5) is Boolean and friendly to encode thus it saves the storage space and boots the performance. In another hand, it is hard-encoded and unfriendly to update dynamically. If we want to train a decision tree(5) from scratch, it requires a more convenient code scheme of the bitvector matrix. Like the Standard Machine in Lebesgue Integral, we generalize the bitvector matrices step by step.
Similar to the extension of activation function, we will obtain the first rule: it does not change the first index of the maximum in $B\sigma(Sx-t)$ for the decision tree (5) if we only replace the number $1$ in the matrix $B$ with the identical positive constants. Observe that $Bh=\sum_{i=1}^{n_{L}}h_{i}B_{i}$, we can positively scale $B_{i}$ with $\alpha_{i}>0$, i.e., $\arg\max(Bh)=\arg\max(B\operatorname{diag}(\alpha)h)$, where $h\in\mathbb{B}^{n_{L}}$ and $B\in\mathbb{R}_{+}^{L\times n_{L}}$ is nonnegative matrix. In another hand, it is equivalent to positively scale $h_{i}$ with $\alpha_{i}>0$.
And when $\arg\max(Bh)=\arg\max(BMh)$, what is the sufficient and necessary condition of the matrix $M$? The matrix $M$ must preserve the first index of maximum in the non-negative vector.
The second rule is to substitute any negative constant for each zero while keep others unchanged over the matrix $B$ in (5), Combining both rules, we would replace the ones with identical positive constants and the zeroes with the negative constants. For example, we can replace the matrix $B$ in 1 with following matrix:
@@
$B_{1}\wedge B_{2}\wedge\cdots B_{m}=B_{1}\circ B_{2}\circ\cdots B_{m}$
@@
$T(x)=v[i],i=\arg\max{[B\,\sigma(Sx-t)+\vec{1}]}$ (7)
where the matrix $B$ is a Boolean/logical matrix, i.e., its element is 00 or $1$; other notations have the same meaning as in (5). This extension makes some sense. For example, if $B$ is the matrix full of ones, the formula (7) is a constant function $T(x)=v[1]=v[2]=\cdots=v[L]$. However, there is no equivalent logical ‘AND’ operation of the columns of the matrix $B$ in above example.
The above representation is the minimal extension of decision trees, where it is just to set the bitvector matrix $B$ free from the four rulesLABEL:four_rules. If the bitvector is extended to real vector space, the vanilla decision tree(5) is sparse in the new representation. If we still keep the bitvector in Boolean space, the classical decision tree is not too rare in our new representation. This extension includes the decision tree. In another word, our new representation is the super set of decision tree.
In (7), the most general form of $B$ is the one whose each element takes the real number as its domain. We call the formula(5) ‘decision machine’ if $B$ is real because most of them cannot be interpreted in the tree-traversal way as the classical original decision trees. We turn to the row space of the bitvector matrix. Each element of the result $Bh$ is the inner product of the row vector of $B$ such as $b$ and the vector $h$, which measures the similarity of the vector $b$ and $h$ in some sense. Thus the result $Bh$ can be interpreted as combination of multiple models, where the largest one is selected to determine the output. And in this case the activation function can extend to negative part of real number. It is less logical, more numerical and closer to neural network.
@@
$T(x)=v[i],i=\arg\max(\tilde{B}\operatorname{sgn}(\tilde{S}\tilde{x}))$ (8)
where $\tilde{B}=\operatorname{diag}(\|B_{1}\|_{1},\cdots,\|B_{L}\|_{1})^{-1}B$; the augmented matrix $\tilde{S}=[S\,\,\vec{b}]\in\mathbb{R}^{m\times(n+1)}$; the augmented vector $\tilde{x}=(x^{t}\,\,-1)^{T}\in\mathbb{R}^{n+1}$; other the notations are same as before except
$B=\begin{pmatrix}-1&-1&0&-1&0\\ -1&-1&0&+1&-1\\ -1&-1&0&+1&+1\\ -1&+1&0&0&0\\ +1&0&-1&0&0\\ +1&0&+1&0&0\\ \end{pmatrix}.$
@@ 3.1.4 The Index Finding Function
$T(x)=\left<\operatorname{softmax}(B\,\sigma(Sx-b)),p(x)\right>=\sum_{i=1}^{L}e_{i}p_{i}(x).$ (9)
Here $(e_{1},\cdots,e_{l})^{T}=\operatorname{softmax}(B\,\sigma(Sx-b))$. In another word we use the softmax function as an alternative to the navigation function in Mohammad Norouzi et al[27].
Note that the we prefer the smaller index of the maximums in the hidden state $h=B\,\sigma(Sx-b)$, it is best to find better approximation of the one-hot vector as shown in [15, 17, 25, 26, 20, 4].
However, all above functions are coordinate independent which does not reflect your preference on minima index of the target leaf node perfectly. Suppose $z\in\mathbb{R}_{+}^{L}$, we would like to find the function that indicate first index of the maximums in the vector $z$. For example, $z=\vec{1}\in\mathbb{R}^{L}$, it is supposed to return
@@
if $z=(1,2,\cdots,L-1,L)^{T}$, it is supposed to return
@@
$\operatorname{sparsemax}(z)=\arg\min_{p\in\triangle_{L-1}}\|p-z\|_{w}^{2}+\|p\|_{1}$
where the norm $\|p-z\|_{w}^{2}=\sum_{i}w_{i}(p_{i}-z_{i})^{2}$. It is a variant of generalized lasso.
@@
All the operators are common in deep learning such as matrix multiplication, ReLU and softmax. There are one affine layer, a nonlinear layer and a softmax layer. In some sense, decision trees are special type neural networks. We call those models ‘logical decision machines’. This representation is ‘mixed-precision’ because only $B$ is 0-1 valued  [^5] In a vanilla binary decision tree, the selection matrix is also 0-1 valued. Here we focus on oblique decision trees. and other notations are real.
@@ 3.1.5 The Prediction
Yu Shi, Jian Li and Zhize Li use piecewise linear regression trees (PL Trees) to enhance the accuracy and efficiency of GBDT. In fact, the prediction functions $p=(p_{1},\cdots,p_{L})^{T}$ can be substituted for the constant output value vector $v=(v_{1},v_{2},\cdots,v_{L})$ associated with the leaves of the decision tree.
$T(x)=p_{i}(x),$ (10)
where $i$ is the first index of maximums of $B\,\sigma(Sx-t)$, i.e., $i=\min\{j\mid v_{j}=\max(v),v=B\,\sigma(Sx-t)\}$; and $p_{i}:\mathbb{R}^{n}\mapsto\mathbb{O}$, $\mathbb{O}$ is the domain of ground truth. other notations are the same as shown in (5). The function $f_{i}$ can be any type function. These are similar to “Multivariate Adaptive Regression Splines”.
@@ 3.2 Unification of Binary Decision Trees
Based on the concept of bitvector matrix $B$, we formulate diverse decision trees in a unified way
$T(x)=\left<p(x),\sigma(h)\right>,h=B\,\tau(f(x))$ (11)
where $i$ is the first index of maximums of the hidden state $h=\tau(f(x))$; $p_{i}(x)$ is the prediction function associated with the leaf node $i$ and $p(x)=(p_{1}(x),\cdots,p_{L}(x))$. other notations are the same as shown in (5). If we still obey the four rules, it is still a decision tree type algorithms. If we set the bitvector free from the four rulesLABEL:four_rules, it is a generalized decision tree with nonlinear test function.
@@ 4 Conclusion and Discussion
We unify diverse decision trees or recursive partitioning methods in our novel parameterized framework of decision trees so that we can use the deep neural networks packages to deploy and optimize the different decision trees. In theory, it is proved that a decision trees is determined by the tuple $(S,t,B,v)$ consisting of 2 matrices and 2 vectors. Specially, we find some mathematical properties of the structure matrix $B$. We present how to unify some extension of decision trees in our new parameterization framework. It is next step to search more efficient methods to jointly optimize the parameters of decision trees. And we will study some theoretical issues in future such as the complexity of our novel representation and the universal approximation of decision trees

### v8

@@ 3.1.3 The Bitvector Matrix
The bitvector matrix $\mathrm{B}$ in (5) is Boolean and friendly to encode thus it saves the storage space and boots the performance. In another hand, it is hard-encoded and unfriendly to update dynamically. If we want to train a decision tree(5) from scratch, it requires a more convenient code scheme of the bitvector matrix. Like the Standard Machine in Lebesgue Integral, we generalize the bitvector matrices step by step.
Similar to the extension of activation function, we will obtain the first rule: it does not change the first index of the maximum in $B\sigma(\mathrm{S}\mathbf{x}-\mathbf{t})$ for the decision tree (5) if we only replace the number $1$ in the matrix $\mathrm{B}$ with the identical positive constants. Observe that $\mathrm{B}\mathbf{b}=\sum_{i=1}^{n_{L}}h_{i}B_{i}$, we can positively scale $B_{i}$ with $\alpha_{i}>0$, i.e., $\arg\max(\mathrm{B}\mathbf{b})=\arg\max(B\operatorname{diag}(\alpha)h)$, where $h\in\mathbb{B}^{n_{L}}$ and $B\in\mathbb{R}_{+}^{L\times n_{L}}$ is nonnegative matrix. In another hand, it is equivalent to positively scale $h_{i}$ with $\alpha_{i}>0$.
And when $\arg\max(\mathrm{B}\mathbf{b})=\arg\max(BMh)$, what is the sufficient and necessary condition of the matrix $M$? The matrix $M$ must preserve the first index of maximum in the non-negative vector.
The second rule is to substitute any negative constant for each zero while keep others unchanged over the matrix $\mathrm{B}$ in (5), Combining both rules, we would replace the ones with identical positive constants and the zeroes with the negative constants. For example, we can replace the matrix $\mathrm{B}$ in 1 with following matrix:
@@
$\mathrm{B}_{1}\wedge\mathrm{B}_{2}\wedge\cdots\mathrm{B}_{m}=\mathrm{B}_{1}\circ\mathrm{B}_{2}\circ\cdots\mathrm{B}_{m}$
@@
$T(\mathbf{x})=v[i],i=\arg\max{[\mathrm{B}\,\sigma(\mathrm{S}\mathbf{x}-\mathbf{t})+\vec{1}]}$ (7)
where the matrix $\mathrm{B}$ is a Boolean/logical matrix, i.e., its element is 00 or $1$; other notations have the same meaning as in (5). This extension makes some sense. For example, if $\mathrm{B}$ is the matrix full of ones, the formula (7) is a constant function $T(\mathbf{x})=v[1]=v[2]=\cdots=v[L]$. However, there is no equivalent logical ‘AND’ operation of the columns of the matrix $\mathrm{B}$ in above example.
If the bitvector is extended to real vector space, the vanilla decision tree(5) is sparse in the new representation. If we still keep the bitvector in Boolean space, the classical decision tree is not too rare in our new representation. This extension includes the decision tree. In another word, our new representation is the super set of decision tree.
In (7), the most general form of $\mathrm{B}$ is the one whose each element takes the real number as its domain. We call the formula(5) ‘decision machine’ if $\mathrm{B}$ is real because most of them cannot be interpreted in the tree-traversal way as the classical original decision trees. We turn to the row space of the bitvector matrix. Each element of the result $\mathrm{B}\mathbf{b}$ is the inner product of the row vector of $\mathrm{B}$ such as $\mathrm{B}$ and the vector $\mathbf{h}$, which measures the similarity of the vector $\mathrm{B}$ and $\mathbf{h}$ in some sense. Thus the result $\mathrm{B}\mathbf{b}$ can be interpreted as combination of multiple models, where the largest one is selected to determine the output. And in this case the activation function can extend to negative part of real number. It is less logical, more numerical and closer to neural network.
@@
$T(\mathbf{x})=v[i],i=\arg\max(\tilde{\mathrm{B}}\operatorname{sgn}(\tilde{S}\tilde{x}))$ (8)
where $\tilde{\mathrm{B}}=\operatorname{diag}(\|B_{1}\|_{1},\cdots,\|B_{L}\|_{1})^{-1}B$; the augmented matrix $\tilde{S}=[S\,\,\vec{b}]\in\mathbb{R}^{m\times(n+1)}$; the augmented vector $\tilde{x}=(x^{t}\,\,-1)^{T}\in\mathbb{R}^{n+1}$; other the notations are same as before except
$\mathrm{B}=\begin{pmatrix}-1&-1&0&-1&0\\ -1&-1&0&+1&-1\\ -1&-1&0&+1&+1\\ -1&+1&0&0&0\\ +1&0&-1&0&0\\ +1&0&+1&0&0\\ \end{pmatrix}.$
@@ 3.1.4 The Index Finding Function
$T(\mathbf{x})=\left<\operatorname{softmax}(\mathrm{B}\,\sigma(\mathrm{S}\mathbf{x}-b)),p(\mathbf{x})\right>=\sum_{i=1}^{L}e_{i}p_{i}(\mathbf{x}).$ (9)
Here $(e_{1},\cdots,e_{l})^{T}=\operatorname{softmax}(\mathrm{B}\,\sigma(\mathrm{S}\mathbf{x}-b))$. In another word we use the softmax function as an alternative to the navigation function in Mohammad Norouzi et al[24].
Note that the we prefer the smaller index of the maximums of $h=\mathrm{B}\,\sigma(\mathrm{S}\mathbf{x}-b)$, it is best to find better approximation of the one-hot vector as shown in [14, 16, 22, 23, 19, 4].
However, all above functions are coordinate independent which does not reflect your preference on minima index of the target leaf node perfectly. Suppose $\mathbf{z}\in\mathbb{R}_{+}^{L}$, we would like to find the function that indicate first index of the maximums in the vector $z$. For example, $\mathbf{z}=\vec{1}\in\mathbb{R}^{L}$, it is supposed to return
@@
if $\mathbf{z}=(1,2,\cdots,L-1,L)^{T}$, it is supposed to return
@@
$\operatorname{sparsemax}(\mathbf{z})=\arg\min_{\mathbf{p}\in\triangle_{L-1}}\|\mathbf{p}-\mathbf{z}\|_{w}^{2}+\|\mathbf{p}\|_{1}$
where the norm $\|\mathbf{p}-\mathbf{z}\|_{w}^{2}=\sum_{i}w_{i}(p_{i}-z_{i})^{2}$. It is a variant of generalized lasso.
@@
All the operators are common in deep learning such as matrix multiplication, ReLU and softmax. There are one affine layer, a nonlinear layer and a softmax layer. In some sense, decision trees are special type neural networks. We call those models ‘logical decision machines’. This representation is ‘mixed-precision’ because only $\mathrm{B}$ is 0-1 valued  [^5] In a vanilla binary decision tree, the selection matrix is also 0-1 valued. Here we focus on oblique decision trees. and other notations are real.
@@ 3.1.5 The Prediction
Yu Shi, Jian Li and Zhize Li use piecewise linear regression trees (PL Trees) to enhance the accuracy and efficiency of GBDT. In fact, the prediction functions $\mathbf{p}=(p_{1},\cdots,p_{L})^{T}$ can be substituted for the constant output value vector $\mathbf{v}=(v_{1},v_{2},\cdots,v_{L})$ associated with the leaves of the decision tree.
$T(\mathbf{x})=p_{i}(\mathbf{x}),$ (10)
where $i$ is the first index of maximums of $\mathrm{B}\,\sigma(\mathrm{S}\mathbf{x}-\mathbf{t})$, i.e., $i=\min\{j\mid v_{j}=\max(v),v=\mathrm{B}\,\sigma(\mathrm{S}\mathbf{x}-\mathbf{t})\}$; and $p_{i}:\mathbb{R}^{n}\mapsto\mathbb{O}$, $\mathbb{O}$ is the domain of ground truth. other notations are the same as shown in (5). The function $f_{i}$ can be any type function. These are similar to “Multivariate Adaptive Regression Splines”.
@@ 3.2 Unification of Binary Decision Trees
Based on the concept of bitvector matrix $\mathrm{B}$, we formulate diverse decision trees in a unified way
$T(\mathbf{x})=\left<p(\mathbf{x}),\sigma(\mathbf{h})\right>,\mathbf{h}=\mathrm{B}\,\tau(f(\mathbf{x}))$ (11)
where $i$ is the first index of maximums of the hidden state $\mathbf{h}=\tau(f(\mathbf{x}))$; $p_{i}(\mathbf{x})$ is the prediction function associated with the leaf node $i$ and $\mathbf{p}=(p_{1}(\mathbf{x}),\cdots,p_{L}(\mathbf{x}))$.
@@ 4 Conclusion and Discussion
We unify diverse decision trees or recursive partitioning methods in our novel parameterized framework of decision trees so that we can use the deep neural networks packages to deploy and optimize the different decision trees. In theory, it is proved that a decision trees is determined by the tuple $(v,\mathrm{B},\mathrm{S},\mathbf{t})$ consisting of 2 matrices and 2 vectors. Specially, we find some mathematical properties of the structure matrix $\mathrm{B}$. We present how to unify some extension of decision trees in our new parameterization framework. It is next step to search more efficient methods to jointly optimize the parameters of decision trees. And we will study some theoretical issues in future such as the complexity of our novel representation and the universal approximation of decision trees